# Spectral Clustering with Likelihood Refinement is Optimal for Latent Class Recovery

### v1

@@ 5.2 Real-data application
We further evaluate the performance of our methods on the United States 112th Senate Roll Call Votes data, which is publicly available at https://legacy.voteview.com/senate112.htm. The dataset contains voting records for $J=486$ roll calls by $102$ U.S. senators. Following the preprocessing procedure in Lyu et al., (2025), we convert the response matrix to binary and remove senators who are neither Democrats nor Republicans, as well as those with more than $10\%$ missing votes. This results in a final sample of $N=94$ senators. For the remaining missing entries, we impute values by randomly assigning 00 or $1$ with probability equal to the positive response rate observed from each senator’s non-missing votes.
@@
[…]
@@ 6 Discussion
Supplementary Material.
The proofs of all theoretical results are included in the Supplementary Material.
Acknowledgement.
@@
Appendix A Supplementary Material
@@ B.1 Proof of Theorem 1
The proof can be adapted from Theorem S.4 in Lyu et al., (2025) by fixing the individual degree parameters therein to be all $1$’s. We only sketch essential modifications here. Let
@@
Following the construction of ${\cal V}^{*}$ the proof of Theorem S.4 in Lyu et al., (2025), it suffices to consider the following space
@@
Then we can follow the proof of Theorem S.4 in Lyu et al., (2025). Notice that in the proof of Lemma S.2, the argument therein still holds by allowing $p_{1,j}\asymp p_{2,j}\asymp 1$ and $\left|p_{1,j}-p_{2,j}\right|=o\left(p_{1,j}\right)$. We can then get the desired result by directly applying the modified Lemma S.2 without further lower bounding it.
@@ B.2 Proof of Theorem 2
In the following, we will derive the clustering error for $i\in{\cal S}_{1}$. Fix any $k\in[K]\backslash\left\{s_{i}^{\star}\right\}$. Without loss of generality, we assume $\pi=\text{Id}$ (identity map with $\pi(k)=k$ for all $k$) in (18) as otherwise we can replace $\theta_{j,k}$ with $\theta_{j,\pi(k)}$ and $\theta_{j,s_{i}^{*}}$ with $\theta_{j,\pi(s_{i}^{*})}$ in the following derivation. By definition we have
@@
where (a) holds due to (18) and $\log\left(1+x\right)\leq x$ for $x>-1$, (b) holds due to $e^{x}-1\leq 2x$ for $10<x<$ 1 and the fact that ${\lambda\rho I^{\star}}/{J}\lesssim{\rho^{1-\epsilon}\Delta^{2}}/{J}=o(1)$, and (c) holds due to the choice of $\lambda=c\rho^{-\varepsilon}$. For the second term of (B.2), by the independence between $\widehat{\theta}^{(2)}_{j,k}$ and $\left\{R_{i,j},j\in[J]\right\}$ we get
@@
^$\left\|\widehat{}\mathbf{\Theta}^{(1)}-\widehat{}\mathbf{\Theta}^{(2)}\mathbf{G}_{\pi_{0}^{-1}}\right\|_{\rm F}^{2}$ $=\sum_{j\in[J]}\sum_{k\in[K]}\left|\widehat{\theta}^{(1)}_{j,k}-\widehat{\theta}_{j,\pi_{0}^{-1}(k)}^{(2)}\right|^{2}$
@@
^$\left\|\widehat{}\mathbf{\Theta}^{(1)}-\widehat{}\mathbf{\Theta}^{(2)}\mathbf{G}_{\pi}\right\|_{\rm F}^{2}$ $=\sum_{j\in[J]}\sum_{k\in[K]}\left|\widehat{\theta}^{(1)}_{j,k}-\widehat{\theta}_{j,\pi(k)}^{(2)}\right|^{2}$
@@ B.3 Proof of Proposition 2
We thus obtain $\lambda_{K}\left(\mathbf{Z}\right)\geq\sqrt{\alpha N/K}$. On the other hand, by Proposition 2 in Lyu et al., (2025) we obtain that for any $\delta\in(0,1)$,
@@
To obtain an lower bound for $\Delta$, it suffices to use Lemma S.1 in Lyu et al., (2025) to get $\Delta\geq\sqrt{2}\lambda_{K}\left(\mathbf{\Theta}\right)$. We get the desired result by setting $\delta=1/2$.
@@ B.4 Proof of Corollary 1
The proof follows the same argument as that of Theorem 2. We only sketch the modifications here. Without loss of generality, we only consider sample data points in ${\cal S}_{1}$. First note that by Proposition 3.1 in Zhang and Zhou, (2024), we have $\ell\left(\widetilde{\bm{s}}^{(2)},{\bm{s}}_{{\cal S}_{2}}^{\star}\right)=o(1)$ with probability at least $1-O\left(e^{-N\vee J}\right)$. Hence we get that
@@ B.5 Proof of Proposition 4
for $0<x<{1}/{2}$. Some algebra gives that
@@
for $0<x<{1}/{2}$, which implies that $g(x)<\lim_{x\rightarrow\frac{1}{2}}g(x)=0$. Hence we have $\tau^{\prime}(x)<0$ and $\tau(x)>\lim_{x\rightarrow\frac{1}{2}}\tau(x)=1$. Thus we have
@@ B.6 Proof of Lemma 1
In light of Corollary 3.12, Remark 3.13 in Bandeira and Van Handel, (2016) and Theorem 6.10 in Boucheron et al., (2013), for $t\geq 0$ we have
@@
This implies that $\lambda_{K+1}\left(\mathbf{R}\right)\leq\left\|\mathbf{E}\right\|\leq 2\left(\sqrt{J}+\sqrt{N}\right)$ with probability $1-o(1)$. On the other hand, by assumption in Proposition 1 on $\sigma_{K}\left(\mathbb{E}\mathbf{R}\right)$, we obtain that
@@
with probability $1-o(1)$. This completes the proof of Lemma 1.
@@
where we’ve used the fact that
$\bar{C}_{\theta+\delta,\theta}=\frac{\theta}{1-\theta}\left(1+o(1)\right).$
This implies that

### v2

@@ 5.2 Real Data Example
We further evaluate the performance of our methods on the United States 112th Senate Roll Call Votes data, which is publicly available at https://legacy.voteview.com/senate112.htm. The dataset contains voting records for $J=486$ roll calls by $102$ U.S. senators. Following the preprocessing procedure in Lyu et al. (2025), we convert the response matrix to binary and remove senators who are neither Democrats nor Republicans, as well as those with more than $10\%$ missing votes. This results in a final sample of $N=94$ senators. For the remaining missing entries, we impute values by randomly assigning 0 or $1$ with probability equal to the positive response rate observed from each senator’s non-missing votes.
@@
We also validate the performance of our method for estimating $K$ on this real dataset. In particular, the threshold defined in (17) is $2.01\times(\sqrt{486}+\sqrt{94})\approx 63.8$. The first three largest singular values of $\mathbf{R}$ are $\{148.1,\,64.4,\,16.6\}$, which leads to an estimated number of clusters $\widehat{K}=2$. Notably, this estimated value exactly matches the known number of clusters in the dataset (i.e., $K=2$).
@@ 6 Discussion
Financial support.
@@
Competing interests.
None.
Appendix A Appendix of the Proofs
@@ B.1 Proof of Theorem 1
The proof can be adapted from Theorem S.4 in Lyu et al. (2025) by fixing the individual degree parameters therein to be all $1$’s. We only sketch essential modifications here. Let
@@
Following the construction of $\cal V$$^{*}$ the proof of Theorem S.4 in Lyu et al. (2025), it suffices to consider the following space
@@
Then we can follow the proof of Theorem S.4 in Lyu et al. (2025). Notice that in the proof of Lemma S.2, the argument therein still holds by allowing $p_{1,j}\asymp p_{2,j}\asymp 1$ and $\left|p_{1,j}-p_{2,j}\right|=o\left(p_{1,j}\right)$. We can then get the desired result by directly applying the modified Lemma S.2 without further lower bounding it.
@@ B.2 Proof of Theorem 2
[…]
@@
where (a) holds due to (18) and $\log\left(1+x\right)\leq x$ for x>1x>-1, (b) holds due to $e^{x}-1\leq 2x$ for $0<x<1$ and the fact that ${\lambda\rho I^{\star}}/{J}\lesssim{\rho^{1-\epsilon}\Delta^{2}}/{J}=o(1)$, and (c) holds due to the choice of $\lambda=c\rho^{-\varepsilon}$. For the second term of (B.2), by the independence between $\widehat{\theta}^{(2)}_{j,k}$ and $\left\{R_{i,j},j\in[J]\right\}$ we get
@@
$\left\|\widehat{}\mathbf{\Theta}^{(1)}-\widehat{}\mathbf{\Theta}^{(2)}\mathbf{G}_{\pi_{0}^{-1}}\right\|_{\rm F}^{2}$ $=\sum_{j\in[J]}\sum_{k\in[K]}\left|\widehat{\theta}^{(1)}_{j,k}-\widehat{\theta}_{j,\pi_{0}^{-1}(k)}^{(2)}\right|^{2}$
@@
$\left\|\widehat{}\mathbf{\Theta}^{(1)}-\widehat{}\mathbf{\Theta}^{(2)}\mathbf{G}_{\pi}\right\|_{\rm F}^{2}$ $=\sum_{j\in[J]}\sum_{k\in[K]}\left|\widehat{\theta}^{(1)}_{j,k}-\widehat{\theta}_{j,\pi(k)}^{(2)}\right|^{2}$
@@ B.3 Proof of Proposition 2
We thus obtain $\lambda_{K}\left(\mathbf{Z}\right)\geq\sqrt{\alpha N/K}$. On the other hand, by Proposition 2 in Lyu et al. (2025) we obtain that for any $\delta\in(0,1)$,
@@
To obtain an lower bound for $\Delta$, it suffices to use Lemma S.1 in Lyu et al. (2025) to get $\Delta\geq\sqrt{2}\lambda_{K}\left(\mathbf{\Theta}\right)$. We get the desired result by setting $\delta=1/2$.
@@ B.4 Proof of Corollary 1
The proof follows the same argument as that of Theorem 2. We only sketch the modifications here. Without loss of generality, we only consider sample data points in ${\cal S}_{1}$. First note that by Proposition 3.1 in Zhang and Zhou (2024), we have $\ell\left(\widetilde{\bm{s}}^{(2)},{\bm{s}}_{{\cal S}_{2}}^{\star}\right)=o(1)$ with probability at least $1-O\left(e^{-N\vee J}\right)$. Hence we get that
@@ B.5 Proof of Proposition 4
for 0<x<1/20<x<{1}/{2}. Some algebra gives that
@@
for 0<x<1/20<x<{1}/{2}, which implies that $g(x)<\lim_{x\rightarrow\frac{1}{2}}g(x)=0$. Hence we have $\tau^{\prime}(x)<0$ and $\tau(x)>\lim_{x\rightarrow\frac{1}{2}}\tau(x)=1$. Thus we have
@@ B.6 Proof of Lemma 1
In light of Corollary 3.12, Remark 3.13 in Bandeira and Van Handel (2016) and Theorem 6.10 in Boucheron et al. (2013), for $t\geq 0$ we have
@@
This implies that $\lambda_{K+1}\left(\mathbf{R}\right)\leq\left\|\mathbf{E}\right\|\leq 2\left(\sqrt{J}+\sqrt{N}\right)$ with probability 1o(1)1-o(1). On the other hand, by assumption in Proposition 1 on $\sigma_{K}\left(\mathbb{E}\mathbf{R}\right)$, we obtain that
@@
with probability 1o(1)1-o(1). This completes the proof of Lemma 1.
@@
where we’ve used the fact that $\bar{C}_{\theta+\delta,\theta}=\frac{\theta}{1-\theta}\left(1+o(1)\right).$ This implies that